# Safety Controller Synthesis for Stochastic Networked Systems under Communication Constraints

### v1

@@ IV Simulation Results
We illustrate the effectiveness of our proposed framework by applying it to an RLC circuit. We consider communication channels with the probabilities of successful data transmission for the uplink and downlink channels given by $p_{\theta}=0.93$ and $q_{\phi}=0.90$, while the delay is $\tau=3$. The RLC circuit dynamics are described as follows [34]
$\begin{cases}x_{1k}=x_{1k}+\rho\left(-\frac{R}{L}x_{1k}-\frac{1}{L}x_{2k}\right)+u_{1k}+w_{1k},\\ x_{2k}=x_{2k}+\rho\left(\frac{1}{C}x_{1k}\right)+u_{2k}+w_{2k},\end{cases}$
where $x_{1k}$ and $x_{2k}$ represent the current and voltage of the circuit, respectively, $\rho=0.05$ is the sampling time, $R=2$ is the series resistance, $L=9$ is the series inductance, $C=0.5$ is the circuit capacitance, and $w_{1k},w_{2k}\sim\mathcal{N}(0,0.1)$.
(a) Open-loop trajectories.
(b) Closed-loop with $\tau=1$.
(c) Closed-loop with $\tau=3$.
(d) Closed-loop with $\tau=4$.
[…]
@@
$\mathcal{F}=\begin{bmatrix}-0.2634&-0.09317\\ -0.09047&-0.2761\end{bmatrix}\!\!.$
With the obtained $\mathcal{P}$ and under Algorithm 2, we proceed to determine the corresponding initial and unsafe level sets $\eta$ and $\beta$, by satisfying conditions (18a) and  (18b). It is worth noting that condition (18b) should be satisfied separately for both unsafe regions (cf. Remark 3). The resulting values are $\eta=0.0226$ and $\beta=0.321$. Furthermore, the constant $c$ is calculated as $9.0927\times 10^{-5}$. Applying Theorem 2, we can guarantee that all trajectories of the RLC circuit with uplink delay of $\tau=3$ initialized from $X_{0}=[-0.4,0.4]^{2}$, remain within the safe set $X\backslash X_{1}$ over the time horizon $\mathcal{T}=100$ with $\mathds{P}\Big{\{}\Sigma\models\Upsilon\Big{\}}\geq 0.9$. As shown in Fig. 2, the control gain $\mathcal{F}$ maintains system safety for uplink delays $\tau\leq 3$. However, the safety specification is violated when the uplink delay $\tau$ reaches $4$ or higher. Fig. 3 displays the number of lost state and control data corresponding to each of the $20$ noise realizations over the time horizon $\mathcal{T}=100$. As an observation, we note that the guaranteed lower bound on the safety probability tends to be conservative, which is expected due to using the CBC as a Lyapunov-like function. In fact, no violations of the safety specification $\Upsilon$ were observed during simulation runs with various noise realizations, for uplink delays $\tau\leq 3$ over the time horizon $\mathcal{T}=100$.
Figure 3: The number of lost state ($\theta_{k}=0$) and lost control ($\phi_{k}=0$) data for RLC circuit with $\tau=3$ over a time horizon of $\mathcal{T}=100$ in the uplink (orange) and downlink (blue) channels for $20$ different simulation runs.
@@ V Conclusion
We proposed a framework for analyzing the safety of dt-SLS experiencing communication imperfections due to wireless network constraints. Specifically, we accounted for constant delays in the uplink channel and data loss modeled as independent Bernoulli processes in both uplink and downlink channels. By introducing an augmented dt-ASLS, we effectively captured the dynamics of the original system under delay and data loss constraints. We employed CBCs to design controllers capable of ensuring safety specifications in a probabilistic sense, translating safety requirements into matrix inequalities. We quantified probabilistic guarantees for maintaining system safety despite communication imperfections. We demonstrated our framework through simulations involving an RLC circuit.
@@ Appendix A Appendix
$\mathcal{A}_{13}$ $=\begin{bmatrix}B\!\!&\!\!\mathbf{0}_{n\times m}\!\!&\!\!\phi AB\!\!&\!\!\dots\!\!&\!\!\phi A^{\tau}B\\ \mathbf{0}_{(\psi-n)\times m}\!\!&\!\!\mathbf{0}_{(\psi-n)\times m}\!\!&\!\!\mathbf{0}_{(\psi-n)\times m}\!\!&\!\!\dots\!\!&\!\!\mathbf{0}_{(\psi-n)\times m}\end{bmatrix}$
$\mathcal{A}_{14}$ $=\begin{bmatrix}B\!\!&\!\!\bar{\phi}AB\!\!&\!\!\dots\!\!&\!\!\bar{\phi}A^{\tau}B\\ \mathbf{0}_{(\psi-n)\times m}\!\!&\!\!\mathbf{0}_{(\psi-n)\times m}\!\!&\!\!\dots\!\!&\!\!\mathbf{0}_{(\psi-n)\times m}\end{bmatrix}_{\psi\times\varpi}$
@@
$\mathcal{A}_{22}$ $=\begin{bmatrix}\boldsymbol{0}_{n\times n}&(1-p_{\theta})A^{2}\\ \mathbf{I}_{n}&\mathbf{0}_{n\times n}\end{bmatrix}_{2n\times 2n}\!\!,\quad\mathcal{A}_{23}=\mathbf{0}_{2n\times(\varpi+m)},$
$\mathcal{A}_{24}$ $=\begin{bmatrix}B&p_{\theta}AB+(1-p_{\theta})B&\dots&p_{\theta}A^{\tau}B\\ \mathbf{0}_{n\times m}&\mathbf{0}_{n\times m}&\dots&\mathbf{0}_{n\times m}\end{bmatrix}_{2n\times\varpi}$
@@
$\mathcal{A}_{32}$ $=\begin{bmatrix}\mathbf{0}_{m\times n}&\phi(1-p_{\theta})\mathcal{F}A\\ \mathbf{0}_{\varpi\times n}&\mathbf{0}_{\varpi\times n}\end{bmatrix}_{(\varpi+m)\times 2n}$
$\mathcal{A}_{33}$ $=\begin{bmatrix}\bar{\phi}\mathbf{I}_{m}&\mathbf{0}_{m\times\varpi}\\ \mathbf{0}_{\varpi\times m}&\mathbf{0}_{\varpi\times\varpi}\end{bmatrix}_{(\varpi+m)\times(\varpi+m)}$
$\mathcal{A}_{34}\!$ $\!=\begin{bmatrix}\phi\mathcal{F}B\!\!&\!\!\phi p_{\theta}\mathcal{F}AB+\phi(1-p_{\theta})\mathcal{F}B\!\!&\!\!\dots\!\!&\!\!\phi p_{\theta}\mathcal{F}A^{\tau}B\\ \mathbf{0}_{\varpi\times m}\!\!&\!\!\mathbf{0}_{\varpi\times m}\!\!&\!\!\dots\!\!&\!\!\mathbf{0}_{\varpi\times m}\end{bmatrix}$
@@
$\mathcal{A}_{42}$ $=\begin{bmatrix}\boldsymbol{0}_{m\times n}&(1-p_{\theta})\mathcal{F}A^{2}\\ \boldsymbol{0}_{(\varpi-m)\times n}&\mathbf{0}_{(\varpi-m)\times n}\end{bmatrix}_{\varpi\times 2n}\mathcal{A}_{43}\!=\!\mathbf{0}_{\varpi\times(\varpi+m)},$
$\mathcal{A}_{44}\!$ $\!=\begin{bmatrix}\mathcal{F}B\!\!&\!\!p_{\theta}\mathcal{F}AB\!+\!(1\!-\!p_{\theta})\mathcal{F}B\!\!&\!\!\dots\!\!&\!\!p_{\theta}\mathcal{F}A^{\tau}B\\ \mathbf{0}_{(\varpi-m)\times m}\!\!&\!\!\mathbf{0}_{(\varpi-m)\times m}\!\!&\!\!\dots\!\!&\!\!\mathbf{0}_{(\varpi-m)\times m}\end{bmatrix}$ (23)
@@
$\mathcal{A}_{14}^{a}$ $=\begin{bmatrix}B\!\!&\!\!\mathbf{0}_{n\times(\varpi-m)}\\ \mathbf{0}_{(\psi-n)\times m}\!\!&\!\!\mathbf{0}_{(\psi-n)\times(\varpi-m)}\end{bmatrix}_{\psi\times\varpi}$
@@
$\mathcal{A}_{14}^{b}$ $=\begin{bmatrix}\mathbf{0}_{n\times m}\!\!&\!\!AB\!\!&\!\!\dots\!\!&\!\!A^{\tau}B\\ \mathbf{0}_{(\psi-n)\times m}\!\!&\!\!\mathbf{0}_{(\psi-n)\times m}\!\!&\!\!\dots\!\!&\!\!\mathbf{0}_{(\psi-n)\times m}\end{bmatrix}_{\psi\times\varpi}$
@@
$\mathcal{A}_{32}^{a}$ $=\begin{bmatrix}\mathbf{0}_{m\times n}&(1-p_{\theta})\mathcal{F}A\\ \mathbf{0}_{\varpi\times n}&\mathbf{0}_{\varpi\times n}\end{bmatrix}_{(\varpi+m)\times 2n}$
$\mathcal{A}_{33}^{a}$ $=\begin{bmatrix}\mathbf{I}_{m}&\mathbf{0}_{m\times\varpi}\\ \mathbf{0}_{\varpi\times m}&\mathbf{0}_{\varpi\times\varpi}\end{bmatrix}_{(\varpi+m)\times(\varpi+m)}$
$\mathcal{A}_{34}^{a}\!$ $\!=\begin{bmatrix}\mathcal{F}B\!\!&\!\!\ p_{\theta}\mathcal{F}AB+(1-p_{\theta})\mathcal{F}B\!\!&\!\!\dots\!\!&\!\!p_{\theta}\mathcal{F}A^{\tau}B\\ \mathbf{0}_{\varpi\times m}\!\!&\!\!\mathbf{0}_{\varpi\times m}\!\!&\!\!\dots\!\!&\!\!\mathbf{0}_{\varpi\times m}\end{bmatrix}$ (24)

### v2

@@ IV Simulation Results
We illustrate the effectiveness of our proposed framework by applying it to an RLC circuit. We consider communication channels with the probabilities of successful data transmission for the uplink and downlink channels given by $p_{\theta}=0.93$ and $q_{\phi}=0.90$, while the delay is $\tau=3$. The RLC circuit dynamics are described as [28]
$\begin{cases}x_{1(k+1)}=x_{1k}+\Delta\left(-\frac{R}{L}x_{1k}-\frac{1}{L}x_{2k}\right)+u_{1k}+w_{1k},\\ x_{2(k+1)}=x_{2k}+\Delta\left(\frac{1}{C}x_{1k}\right)+u_{2k}+w_{2k},\end{cases}$
where $x_{1k}$ and $x_{2k}$ represent the current and voltage of the circuit, respectively, $\Delta=0.05$ is the sampling time, $R=2$ is the series resistance, $L=9$ is the series inductance, $C=0.5$ is the circuit capacitance, and $w_{1k},w_{2k}\sim\mathcal{N}(0,0.1)$.
(a) Open-loop trajectories
(b) Closed loop with $\tau=1$
(c) Closed loop with $\tau=3$
(d) Closed loop with $\tau=4$
Figure 2: Trajectories of the RLC circuit over 20 simulation runs with distinct noise realizations starting from the initial set $X_{0}=\left[-0.4,\,0.4\right]^{2}$. Initial and unsafe regions are depicted by green  and red  boxes, respectively. Plot (a) shows the open-loop trajectories, while the remaining subplots show closed-loop trajectories with the proposed controller $\mathcal{F}$ with delays of $\tau=1$ (b) and $\tau=3$ (c). Plot (d) illustrates trajectories with delay $\tau=4$ over $\mathcal{T}=100$, with the controller gain (22) designed for $\tau\leq 3$, showing violation of the safety property $\Upsilon$, as expected.
@@
$\mathcal{F}=\begin{bmatrix}-0.2634&-0.09317\\ -0.09047&-0.2761\end{bmatrix}\!\!$. (22)
With the obtained $\mathcal{P}$ and under Algorithm 2, we proceed to determine the corresponding initial and unsafe level sets $\eta$ and $\beta$, by satisfying conditions (17a) and  (17b), as $\eta=0.0228$ and $\beta=0.3131$. It is worth noting that condition (17b) should be satisfied separately for both unsafe regions. Furthermore, the constant $c$ is calculated as $7.6307\times 10^{-5}$. Applying Theorem 2, we can guarantee that all trajectories of the RLC circuit with uplink delay of $\tau=3$ initialized from $X_{0}=[-0.4,0.4]^{2}$, remain within the safe set $X\backslash X_{1}$ over the time horizon $\mathcal{T}=100$ with $\mathds{P}\Big{\{}\Sigma\models\Upsilon\Big{\}}\geq 0.9$.
As shown in Fig. 2, the control gain $\mathcal{F}$ maintains system safety for uplink delays $\tau\leq 3$. In fact, no violations of the safety specification $\Upsilon$ were observed during simulation runs with various noise realizations, for uplink delays $\tau\leq 3$ over the time horizon $\mathcal{T}=100$. However, the safety specification is violated when the uplink delay $\tau$ reaches $4$ or higher. Fig. 3 displays the number of lost state and control data corresponding to each of the $20$ noise realizations over the time horizon $\mathcal{T}=100$.
It is worth noting that the controller in (22) is designed for a delay of $\tau=3$ and therefore cannot satisfy the system’s safety specification when subjected to delays $\tau>3$, as illustrated in Fig. 2, subplot (d). Nevertheless, for delays $\tau>3$ and different values of $q_{\phi}$ and $p_{\theta}$, one may attempt to re-solve the condition in (14) together with those in (17) to synthesize safety controllers and their corresponding barrier certificates. We also note that the role of the downlink transmission success probability, $q_{\phi}$, in condition (14) shows that increasing $q_{\phi}$ improves the likelihood of satisfying condition (14), as expected. A similar effect is anticipated for the uplink transmission success probability, $p_{\theta}$.
Figure 3: The number of lost states ($\theta_{k}=0$) and lost control ($\phi_{k}=0$) data for RLC circuit with $\tau=3$ over a time horizon of $\mathcal{T}=100$ in the uplink (orange) and downlink (blue) channels for 2020 different simulation runs.
@@ V Conclusion
We proposed a framework for analyzing the safety of dt-SLS experiencing communication imperfections due to wireless network constraints. Specifically, we accounted for constant delays in the uplink channel and data loss modeled as independent Bernoulli processes in both uplink and downlink channels. By introducing an augmented dt-ASLS, we effectively captured the dynamics of the original system under delay and data loss constraints. We employed CBCs to design controllers capable of ensuring safety specifications in a probabilistic sense, translating safety requirements into matrix inequalities. We quantified probabilistic guarantees for maintaining system safety despite communication imperfections. We demonstrated our framework through simulations involving an RLC circuit. Extending our framework to accommodate networked control systems with nonlinear dynamics is under investigation as future work.
@@ Appendix A Appendix
$\mathcal{A}_{13}$ $=\begin{bmatrix}B\!\!&\!\!\mathbf{0}_{n\times m}\!\!&\!\!\bar{\phi}AB\!\!&\!\!\dots\!\!&\!\!\bar{\phi}A^{\tau}B\\ \mathbf{0}_{(\psi-n)\times m}\!\!&\!\!\mathbf{0}_{(\psi-n)\times m}\!\!&\!\!\mathbf{0}_{(\psi-n)\times m}\!\!&\!\!\dots\!\!&\!\!\mathbf{0}_{(\psi-n)\times m}\end{bmatrix}$
$\mathcal{A}_{14}$ $=\begin{bmatrix}\mathbf{0}_{n\times m}\!\!&\!\!\phi AB\!\!&\!\!\dots\!\!&\!\!\phi A^{\tau}B\\ \mathbf{0}_{(\psi-n)\times m}\!\!&\!\!\mathbf{0}_{(\psi-n)\times m}\!\!&\!\!\dots\!\!&\!\!\mathbf{0}_{(\psi-n)\times m}\end{bmatrix}_{\psi\times\varpi}$
@@
$\mathcal{A}_{22}$ $=\begin{bmatrix}\boldsymbol{0}_{n\times n}&(1-p_{\theta})A^{2}\\ \mathbf{I}_{n}&\mathbf{0}_{n\times n}\end{bmatrix}_{2n\times 2n}\!\!,\quad\mathcal{A}_{23}=\mathbf{0}_{2n\times(\varpi+m)}$
$\mathcal{A}_{24}$ $=\begin{bmatrix}B&AB&p_{\theta}A^{2}B&\dots&p_{\theta}A^{\tau}B\\ \mathbf{0}_{n\times m}&\mathbf{0}_{n\times m}&\mathbf{0}_{n\times m}&\dots&\mathbf{0}_{n\times m}\end{bmatrix}_{2n\times\varpi}$
@@
$\mathcal{A}_{32}$ $=\begin{bmatrix}\mathbf{0}_{m\times n}&\phi(1-p_{\theta})\mathcal{F}A^{2}\\ \mathbf{0}_{\varpi\times n}&\mathbf{0}_{\varpi\times n}\end{bmatrix}_{(\varpi+m)\times 2n}$
$\mathcal{A}_{33}$ $=\begin{bmatrix}\bar{\phi}\begin{bmatrix}\mathbf{I}_{m}&\mathbf{0}_{m\times(\varpi-m)}\end{bmatrix}_{m\times\varpi}&\mathbf{0}_{m\times m}\\ \mathbf{I}_{\varpi\times\varpi}&\mathbf{0}_{\varpi\times m}\end{bmatrix}_{(\varpi+m)\times(\varpi+m)}$
$\mathcal{A}_{34}\!$ $\!=\begin{bmatrix}\phi\mathcal{F}B\!\!&\!\!\phi\mathcal{F}AB\!\!&\!\!\phi p_{\theta}\mathcal{F}A^{2}B\!\!&\!\!\dots\!\!&\!\!\phi p_{\theta}\mathcal{F}A^{\tau}B\\ \mathbf{0}_{\varpi\times m}\!\!&\!\!\mathbf{0}_{\varpi\times m}\!\!&\!\!\mathbf{0}_{\varpi\times m}\!\!&\!\!\dots\!\!&\!\!\mathbf{0}_{\varpi\times m}\end{bmatrix}$
@@
$\mathcal{A}_{42}$ $=\begin{bmatrix}\boldsymbol{0}_{m\times n}&(1-p_{\theta})\mathcal{F}A^{2}\\ \boldsymbol{0}_{(\varpi-m)\times n}&\mathbf{0}_{(\varpi-m)\times n}\end{bmatrix}_{\varpi\times 2n}\!\!,\quad\mathcal{A}_{43}\!=\!\mathbf{0}_{\varpi\times(\varpi+m)}$
$\mathcal{A}_{44}\!$ $\!=\begin{bmatrix}\begin{bmatrix}\mathcal{F}B\!\!&\!\!\mathcal{F}AB\!\!&\!\!p_{\theta}\mathcal{F}A^{2}B\!\!&\!\!\dots\!\!&\!\!p_{\theta}\mathcal{F}A^{(\tau-1)}B\end{bmatrix}\!\!&\!\!p_{\theta}\mathcal{F}A^{\tau}B\\ \mathbf{I}_{(\varpi-m)\times(\varpi-m)}\!\!&\!\!\mathbf{0}_{(\varpi-m)\times m}\end{bmatrix}$ (23)
@@
$\mathcal{A}_{14}^{a}$ $=\begin{bmatrix}\mathbf{0}_{n\times m}\!\!&\!\!AB\!\!&\!\!\dots\!\!&\!\!A^{\tau}B\\ \mathbf{0}_{(\psi-n)\times m}\!\!&\!\!\mathbf{0}_{(\psi-n)\times m}\!\!&\!\!\dots\!\!&\!\!\mathbf{0}_{(\psi-n)\times m}\end{bmatrix}_{\psi\times\varpi}$
@@
$\mathcal{A}_{32}^{a}$ $=\begin{bmatrix}\mathbf{0}_{m\times n}&(1-p_{\theta})\mathcal{F}A^{2}\\ \mathbf{0}_{\varpi\times n}&\mathbf{0}_{\varpi\times n}\end{bmatrix}_{(\varpi+m)\times 2n}$
$\mathcal{A}^{a}_{33}$ $=\begin{bmatrix}\mathbf{0}_{m\times\varpi}&\mathbf{0}_{m\times m}\\ \mathbf{I}_{\varpi\times\varpi}&\mathbf{0}_{\varpi\times m}\end{bmatrix}_{(\varpi+m)\times(\varpi+m)}$
$\mathcal{A}^{b}_{33}$ $=\begin{bmatrix}\begin{bmatrix}\mathbf{I}_{m}&\mathbf{0}_{m\times(\varpi-m)}\end{bmatrix}_{m\times\varpi}&\mathbf{0}_{m\times m}\\ \mathbf{0}_{\varpi\times\varpi}&\mathbf{0}_{\varpi\times m}\end{bmatrix}_{(\varpi+m)\times(\varpi+m)}$
$\mathcal{A}_{34}^{a}\!$ $\!=\begin{bmatrix}\mathcal{F}B\!\!&\!\!\mathcal{F}AB\!\!&\!\!p_{\theta}\mathcal{F}A^{2}B\!\!&\!\!\dots\!\!&\!\!p_{\theta}\mathcal{F}A^{\tau}B\\ \mathbf{0}_{\varpi\times m}\!\!&\!\!\mathbf{0}_{\varpi\times m}\!\!&\!\!\mathbf{0}_{\varpi\times m}\!\!&\!\!\dots\!\!&\!\!\mathbf{0}_{\varpi\times m}\end{bmatrix}$ (24)